OpenStudy (anonymous):

3-3sinx-2cos^2x = 0 Find all solutions in the domain [-pi,pi]

OpenStudy (anonymous):

Just change cos2x into cos^2x-sin^2x and solve quadratic equation

OpenStudy (mertsj):

\[3-3\sin x-2(1-\sin ^2x)=0\]

OpenStudy (anonymous):

I just need someone to check answers with I got \[-\pi/2, \pi/2, 2\pi/3, -2\pi/3\]

OpenStudy (mertsj):

\[3-3\sin x-2+2\sin ^2x=0\]

OpenStudy (mertsj):

Not what I got.

OpenStudy (anonymous):

wait I mean

OpenStudy (anonymous):

\[-\pi/2 , \pi/2 , 5\pi/6, -5\pi/6 , -\pi/6, \pi/6\]

OpenStudy (mertsj):

Not what I got.

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

Once I factored the quadratic I got (2sinx-1) (sinx-1)=0

OpenStudy (mertsj):

\[\frac{\pi}{6}, \frac{5\pi}{6}, \frac{\pi}{2}\]

OpenStudy (mertsj):

Yes. So sinx=1/2 or 1

OpenStudy (anonymous):

The domain is [-pi,pi] though?

OpenStudy (mertsj):

Which is the same thing as [0,2pi]

OpenStudy (anonymous):

Wait I don't get how that's the same?

OpenStudy (anonymous):

The domain is negative pi to pi? How is it 0 to 2pi?

OpenStudy (mertsj):

That means give all the solutions between - pi and pi . -pi to 0 would include quadrants 3 and 4. 0 to pi would include quadrants 1 and 2

OpenStudy (anonymous):

But -pi/2 and -5pi/6 is between them?

OpenStudy (mertsj):

The sin of -pi/2 is -1 The sin of -5pi/6 is -1/2

OpenStudy (mertsj):

So those cannot be solutions.

OpenStudy (mertsj):

And besides -pi/2 is the same as 3pi/2

OpenStudy (mertsj):

And -5pi/5 is the same as 7pi/6

OpenStudy (mertsj):

-5pi/6. Sorry for the typo

OpenStudy (anonymous):

sin of -1 is 3pi/2 though

OpenStudy (mertsj):

Do you mean that the sin of 3pi/2 is -1?

OpenStudy (anonymous):

yeah, you said "The sin of -pi/2 is -1"

OpenStudy (anonymous):

How can they both be?

OpenStudy (mertsj):

And that is correct because the sin of -pi/2 is the same as the sin of 3pi/2

OpenStudy (mertsj):

Let me draw you a picture.

OpenStudy (mertsj):